OpenStudy (anonymous):

How do you find the exact values for the six trigonometric funcitons of the angle theta, when it's in standard posistion, and you're given the coordinates (-8/17, 15/17)?

OpenStudy (anonymous):

Standard position = quadrant 1 Therefore x = 15, y = 17 tan theta = opp/adj theta = arctan (17/15)

OpenStudy (anonymous):

You need to create a triangle where the angle is in standard position. When you set up the triangle, the signs should all become positive. Then using the triangle, you find sin=opp/hyp cos=adj/hyp tan=opp/adj csc=hyp/opp sec=hyp/adj cot=adj/opp Then you change the values to be positive or negative depending on the quadrant it is in

OpenStudy (anonymous):

what do I do with the -8/17 though? I just don't understand how to make a triangle, because it gives us a circle, with the other side of the triangle in the 2nd quadrant..

OpenStudy (anonymous):

Could you draw out what you mean?

OpenStudy (anonymous):